Math 131 - section 01 - Precalculus Diagnostic Quiz August 31, 2007 Your Name: Circle the correct answer. Show your work. Please turn over for problems 4,5,6.

1. Simplify: \sqrt{15} (\frac{1}{\sqrt{3}} + \frac{1}{\sqrt{5}})

A. 2\begin{matrix}√ & \overline{15} \\ √ & \overline{8} \end{matrix} B. \frac{\sqrt{15}}{\sqrt{3} + \sqrt{5}} C. 1 D. \sqrt{3} + \sqrt{5} E. 2\sqrt{3}

2. Simplify: \frac{12x}{3x − 6} · \frac{x^{2} − 4}{2x + 4}

A. \frac{x^{2} + 12x − 4}{5x − 2}  B. 2x C. −\frac{4}{3}(x − 1) D. \frac{2x(x − 2)}{x + 2}  E. \frac{12x^{3} − 48x}{6x^{2} − 12}

3. Simplify: (u^{−5}v^{2})^{3} (\frac{v^{2}}{u} )^{−1}

A. u^{−16}v^{8} B. u^{−14}v^{4} C. u^{−9}v^{3} D. (uv)^{−7} E. u^{−4}v^{2} 4. If f(x) = 5x^{2} − 11, what is f(a − 2)?

5. Find all values of x satisfying 2(x − 2) > 5.

6. Solve for x: 2x^{2} − x − 6 = 0 (find all solutions).